OpenStudy (anonymous):

The loudness of a sound in decibels is related to the intensity of the sound i, where the function for the loudness is given by L(i)=101og10(i/i0), and io is a constant equal to the minimum sound that can be detected by the human ear. a. Find the loudness in decibels of a rock music recording with an intensity of 10 million times io.

OpenStudy (asnaseer):

\[L(i)=10\log_{10}(\frac{i}{i_0})\]and you are told to find loudness of music that has:\[i=10,000,000*i_0=10^7*i_0\]so you get:\[L(i)=10\log_{10}(\frac{10^7*i_0}{i_0})\]the \(i_0\)'s cancel and you should then be able to calculate the loudness.

OpenStudy (anonymous):

So it would come out to be L(i)=10log10(10^7)?

OpenStudy (anonymous):

Then I calculate?

OpenStudy (asnaseer):

yes, and you should be able to simplify the log

OpenStudy (asnaseer):

\[\log_{10}(10^a)=a\]

OpenStudy (asnaseer):

therefore:\[\log_{10}(10^7)=?\]

OpenStudy (anonymous):

7? http://www.wolframalpha.com/input/?i=log10%2810%5E7%29

OpenStudy (asnaseer):

yes 7, but you didn't need wolfram for that, I showed you the rule above ^^^

OpenStudy (asnaseer):

so now you should be able to calculate the value for loudness

OpenStudy (anonymous):

Oh woops :/ Yea, thank you! a refers to the 7.

OpenStudy (asnaseer):

yes :)

OpenStudy (anonymous):

So would the answer be 7 million?

OpenStudy (asnaseer):

no....

OpenStudy (asnaseer):

\[L(i)=10\log_{10}(\frac{10^7*i_0}{i_0})=10\log_{10}(10^7)=10*7=?\]

OpenStudy (anonymous):

I meant 70 million. But nvm. 10*7=70

OpenStudy (asnaseer):

yes, answer is 70 (not 70 million)

OpenStudy (anonymous):

Oh o.k. Thank you so much for your help asnaseer!

OpenStudy (asnaseer):

yw - I'm glad I was able to help :)

OpenStudy (anonymous):

Would you happen to understand this question? c.In terms of decibels, how close is the intensity level of the rock music recording to the threshold of pain?

OpenStudy (asnaseer):

I think you'll need to google that. try and search for how many decibels a human ear can stand before it becomes painful. for rock music I guess the loudness is 70 decibels from the equation above.

OpenStudy (anonymous):

O.k cool. Thanks!

OpenStudy (anonymous):

IIRC the threshold of pain for the human ear is 120dB, but you might wanna check that